### v1

@@ 3. $(\mathcal{S}_{3})$
From Remark 3.10, it is sufficient to show that for each $\mu\in F$, there exists a facet $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$.
Let $\mu\in F$ and $F^{\prime c}:=(F^{c}\setminus\{n-1\})\sqcup\{\mu\}=\{\alpha_{s},s_{1},\mu\}$. Since $\alpha_{s}\leq m-2$, we have $32m-\alpha_{s}-1\geq m+1\geq\alpha_{s}+$ 3, and hence $32m-\alpha_{s}-1\geq\alpha_{s}+$ 3. First, suppose that $s_{1}\neq\alpha_{s}+4$. Then $s_{1}\in\mathcal{U}$ implies that $s_{1}\notin\{\alpha_{s}-4\text{ (mod $n$)},\alpha_{s}-3,\alpha_{s}-2,\ldots,\alpha_{s}+4\}\cup\{n-1\}$. This means that $\alpha_{s}\nsim s_{1}$ and $|\alpha_{s}-s_{1}|\geq 5$. Further, since $\alpha_{s}\geq 3$ and $s_{1}<n-1$, we get $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. Thus $F^{\prime}\in M(\Delta_{3}(W_{n}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime}$. So, let $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime}\ll F$ by Proposition 3.11. Since $\alpha_{s}<m-1$, we have $F,F^{\prime}\notin\mathcal{D}$. Hence, $F^{\prime}\prec F$ by Definition 3.4 (i). Thus, we take $\tilde{F}=F^{\prime}$.
Now suppose that $s_{1}=\alpha_{s}+4$. First, let $\mu\neq\alpha_{s}+2$. We have $13\leq\alpha_{s}<\alpha_{s}+4=s_{1}<n-$ 1. It follows that $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. Further, since $\alpha_{s}\nsim s_{1}$, we get $F^{\prime}\in M(\Delta_{3}(W_{n}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime}$. So assume that $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime}\ll F$ by Proposition 3.11. Since $\alpha_{s}<m-1$, we have $F,F^{\prime}\notin\mathcal{D}$. Therefore, $F^{\prime}\prec F$ by Definition 3.4 (i), and we take $\tilde{F}=F^{\prime}$.
Now, let $\mu=\alpha_{s}+2$. Let $F^{\prime\prime c}:=(F^{c}\setminus\{s_{1}\})\sqcup\{\mu\}=\{\alpha_{s},\mu,n-1\}$. Since $13\leq\alpha_{s}<\alpha_{s}+2=\mu=s_{1}-2<s_{1}<n-$ 1, it follows that $\alpha_{s}\nsim n-1$ and $\mu\nsim n-1$. Therefore, $W_{n}[F^{\prime\prime c}]$ is disconnected. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime\prime}\prec F$ by Proposition 3.5 and we take $\tilde{F}=F^{\prime\prime}$. So, let $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime\prime}\in\mathcal{T}_{s}$ and thus, $F^{\prime\prime c}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. Since $\mu=\alpha_{s}+2<\alpha_{s}+4=s_{1}$, we have $F^{\prime\prime}\ll F$. Also, $\alpha_{s}<m-1$ implies that $F,F^{\prime\prime}\notin\mathcal{D}$. Hence, $F^{\prime\prime}\prec F$ by Definition 3.4 (i) and we take $\tilde{F}=F^{\prime\prime}$.
Therefore, in each case, we have a facet $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$. Thus, $F$ is a spanning facet. ∎
@@
Since $F\in\mathcal{S}$, we have $F\in\mathcal{T}_{s}$ for some $s$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$. Let $\mathcal{U}=V(W_{n})\setminus\bigl{(}\{\alpha_{s}-4\ (\text{mod $n$}),\alpha_{s}-3,\alpha_{s}-2,\ldots,2m-\alpha_{s}-1\}\cup\{n-1\}\bigr{)}$.
@@
    We have $\alpha_{s}=3$ and $s_{1}\in V(W_{n})\setminus(\{0,1,2,3,\ldots,2m-4\}\sqcup\{n-1\})$. Observe that $s_{1}\in\mathcal{U}$. Therefore, $F$ is a spanning facet by Proposition 3.12.
@@
    We have $\alpha_{s}\in\{4,5,\ldots,m-2\}$ and $s_{1}\in V(W_{n})\setminus\{\alpha_{s}-4,\alpha_{s}-3,\alpha_{s}-2\}\sqcup\{\alpha_{s},\alpha_{s}+1,\ldots,2m-\alpha_{s}-1\}\sqcup\{n-1\}$.
@@
    So assume that $s_{1}=\alpha_{s}-1$. We consider three cases: $\mu<\alpha_{s}$, $\mu\in\{\alpha_{s}+1,\alpha_{s}+2\}$, and $\mu>\alpha_{s}+2$.
    First, let $\mu<\alpha_{s}$. Then $\mu<s_{1}=\alpha_{s}-1$ (as $\mu\in F$). Moreover, $\alpha_{s}\leq m-2$ implies that $\alpha_{s}<_{\mathcal{O}}\mu$. Hence $F^{\prime\prime}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. Since $64\leq\alpha_{s}\leq m-2\leq n-$ 6, we have $\alpha_{s}\nsim n-1$. Further, $\mu<\alpha_{s}$ implies that either $\mu\in\{0,1\}$, or $2\leq\mu<\alpha_{s}$. Hence, $\mu\nsim\alpha_{s}$ or $\mu\nsim n-1$. Therefore $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $\mu<s_{1}$, we have $F^{\prime\prime}\ll F$. Observe that $F,F^{\prime\prime}\notin\mathcal{D}$ (as $\alpha_{s}<m-1$). Therefore, $F^{\prime\prime}\prec F$ Definition 3.4 (i) and we take $\tilde{F}=F^{\prime\prime}$.
    Now, let $\mu\in\{\alpha_{s}+1,\alpha_{s}+2\}$. We have $14\leq\alpha_{s}<\mu\leq\alpha_{s}+2\leq m\leq n-4<n-$ 1. Thus, $\alpha_{s}\nsim n-1$ and $\mu\nsim n-1$. Hence $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $\alpha_{s}\leq m-2$, we have $\mu<_{\mathcal{O}}\alpha_{s}$. Therefore, $F^{\prime\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime\prime}$.
    Lastly, assume that $\mu>\alpha_{s}+2$. Then $13\leq\alpha_{s}-1=s_{1}<\alpha_{s}<\alpha_{s}+2<\mu<n-$ 1 implies that $\mu\nsim\alpha_{s}$ and $\mu\nsim s_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime}$. Now, if $\alpha_{s}<_{\mathcal{O}}\mu$, then $F^{\prime}\ll F$ by Proposition 3.11. Since $\alpha_{s}<m-1$, we have $F,F^{\prime}\notin\mathcal{D}$. Hence, $F^{\prime}\prec F$ by Definition 3.4 (i). Therefore, we take $\tilde{F}=F^{\prime}$.
@@
    We have $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}\in\{m-1,m,m+1,m+2,\ldots,n-3\}$, $s_{1}\in V(W_{n})\setminus(\{\omega,\omega+1,\ldots,\alpha_{s},\alpha_{s}+1,\alpha_{s}+2,\alpha_{s}+3\text{ (mod $n$})\}\sqcup\{y\})$, $\omega=\min\{2m-\alpha_{s},\alpha_{s}-4\}$ and
@@
    First, suppose that $s_{1}\neq\alpha_{s}+4$ (mod $n$). Since $s_{1}\notin\{\omega,\omega+1,\ldots,\alpha_{s}+3\text{ (mod $n$)}\}$ and $\omega\leq\alpha_{s}-4$, it follows that $s_{1}<\alpha_{s}-4$ or $s_{1}>\alpha_{s}+4$ (mod $n$). Thus $|\alpha_{s}-s_{1}|\geq 5$. Now, since $\alpha_{s}\geq m-1>3$ and $s_{1}<n-1$, we get $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. We have $\alpha_{s}\nsim s_{1}$. Therefore $F^{\prime}\in M(\Delta_{3}(W_{n}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime}$. Now, if $\alpha_{s}<_{\mathcal{O}}\mu$, then $F^{\prime}\ll F$ by Proposition 3.11. Since $s_{1}<\alpha_{s}-4$ or $s_{1}>\alpha_{s}+4$ (mod $n$), we have $F,F^{\prime}\notin\mathcal{D}$. Hence $F^{\prime}\prec F$ by Definition 3.4 (i). Therefore, we take $\tilde{F}=F^{\prime}$.
    Now, let $s_{1}=\alpha_{s}+4$ (mod $n$). Observe that if $\alpha_{s}\geq n-5$, then $\alpha_{s}+4\ \text{(mod $n$)}=y$. Therefore, since $s_{1}\neq y$, we have $\alpha_{s}\leq n-6$. This means that $\alpha_{s}+4\ \text{(mod $n$)}=\alpha_{s}+4$. First, let $\mu\notin\{\alpha_{s}+2,\alpha_{s}+3\}$. We have $14\leq m-1\leq\alpha_{s}<\alpha_{s}+4=s_{1}<n-$ 1. It follows that $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. Further, since $\alpha_{s}\nsim s_{1}$, we get $F^{\prime}\in M(\Delta_{3}(W_{n}))$.
    If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.5, and we take $\tilde{F}=F^{\prime}$. So assume that $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime}\ll F$ by Proposition 3.11. Since $n-1>s_{1}=\alpha_{s}+4$ and $\mu\neq\alpha_{s}+3$, we have $F,F^{\prime}\notin\mathcal{D}$. Therefore, $F^{\prime}\prec F$ by Definition 3.4 (i), and we take $\tilde{F}=F^{\prime}$.
    Now, let $\mu\in\{\alpha_{s}+2,\alpha_{s}+3\}$. Observe that since $\alpha_{s}\geq m-1$, we have $\alpha_{s}<_{\mathcal{O}}\mu$, and hence $F^{\prime\prime}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. We have $14\leq m-1\leq\alpha_{s}<\alpha_{s}+2<\alpha_{s}+3\leq n-3<n-$ 1 (as $\alpha_{s}\leq n-6$). Clearly, $\alpha_{s}\nsim n-1$. Moreover, if $\mu=\alpha_{s}+2$, then $\mu\nsim n-1$; and if $\mu=\alpha_{s}+3$, then $\mu\nsim\alpha_{s}$. Thus $F^{\prime\prime}\in M(\Delta_{3}(W_{n}))$. Since $\mu<s_{1}$, we have $F^{\prime\prime}\ll F$. Further, $\alpha_{s}\leq n-6$ implies that $n-1\geq\alpha_{s}+5$ and thus, $F,F^{\prime\prime}\notin\mathcal{D}$. Hence, $F^{\prime\prime}\prec F$ by Definition 3.4 (i), and we take $\tilde{F}=F^{\prime\prime}$.
    Thus, in each case, we have a facet $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$. Thus, $F$ is a spanning facet.
@@
Suppose $F\in M(\Delta_{3}(W_{n}))$ be such that either $F^{c}=\{x,x+1\ (\text{mod $n$}),x+4\ (\text{mod $n$})\}$, or $F^{c}=\{x,x+3\ (\text{mod $n$}),x+4\ (\text{mod $n$})\}$ for some $x\in V(W_{n})$. Then $F$ is not a spanning facet.
@@
We have $x+2$ (mod $n$) $\in F$. Let ${F^{\nu}}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{x+2\text{ (mod $n$)}\}$ for $\nu\in F^{c}$. Then for any $\nu\in F^{c}$, $W_{n}[{F^{\nu}}^{c}]$ is connected. Hence, there does not exist a facet $\tilde{F}$ such that $\tilde{F}\prec F$ and $\tilde{F}\cap F=F\setminus\{x+2\text{ (mod \ $n$)}\}$. Therefore, $F$ is not a spanning facet by definition. ∎
@@
Observe that if $F\in\mathcal{D}$, then either $F^{c}=\{x,x+1\ (\text{mod \ $n$}),x+4\ (\text{mod \ $n$})\}$ or $F^{c}=\{x,x+3\ (\text{mod \ $n$}),x+4\ (\text{mod \ $n$})\}$ for some $x\in V(W_{n})$. Therefore, if $F\in\mathcal{D}$, then $F$ is not a spanning facet by Proposition 3.14.
@@
Let $F,F^{\prime}\in M(\Delta_{3}(W_{n}))$. Let $F\in\mathcal{T}_{s}$ and $F^{\prime}\in\mathcal{T}_{t}$ for some $s\neq t$ such that ${F}^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$ and ${F^{\prime}}^{c}=\{\mu,s_{1},s_{2}\}$ for some $\mu\in F$. If $\alpha_{s}<_{\mathcal{O}}\mu$, then $F\prec F^{\prime}$.
@@
Let $F\in M(\Delta_{3}(W_{n}))$ be a spanning facet such that $F\in\mathcal{T}_{s}$ for some $s$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$. Then $\alpha_{s}\notin\{0,1,2,n-2,n-1\}$ and $s_{2}=n-1$.
@@
By Remark 3.3, we have $\alpha_{s}\notin\{0,1,n-1\}$. Now, since $F$ is a spanning facet, we get $\alpha_{s}\notin\{2,n-2\}$ by Proposition 3.14. Hence $\alpha_{s}\notin\{0,1,2,n-2,n-1\}$.
Observe that $\alpha_{s}<_{\mathcal{O}}n-1$. Suppose $n-1\notin F^{c}$. Then Remark 3.10 implies that there exists $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{n-1\}$ for some $\nu\in F^{c}$.

### v2

@@ 3. $(\mathcal{S}_{3})$
From Remark 3.11, it is sufficient to show that for each $\mu\in F$, there exists a facet $\tilde{F}\in M(\Delta_{3}(C_{n}^{2}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$.
Let $\mu\in F$ and $F^{\prime c}:=(F^{c}\setminus\{n-1\})\sqcup\{\mu\}=\{\alpha_{s},s_{1},\mu\}$. Since $\alpha_{s}\leq m-2$, we have $32m-\alpha_{s}-1\geq m+1\geq\alpha_{s}+$ 3, and hence $32m-\alpha_{s}-1\geq\alpha_{s}+$ 3. First, suppose that $s_{1}\neq\alpha_{s}+4$. Then $s_{1}\in\mathcal{U}$ implies that $s_{1}\notin\{\alpha_{s}-4\text{ (mod $n$)},\alpha_{s}-3,\alpha_{s}-2,\ldots,\alpha_{s}+4\}\cup\{n-1\}$. This means that $\alpha_{s}\nsim s_{1}$ and $|\alpha_{s}-s_{1}|\geq 5$. Further, since $\alpha_{s}\geq 3$ and $s_{1}<n-1$, we get $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. Thus $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.6, and we take $\tilde{F}=F^{\prime}$. So, let $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime}\in\mathcal{T}_{s}$ and $F^{\prime}\ll F$ by Proposition 3.12. Since $\alpha_{s}<m-1$, we have $F,F^{\prime}\notin\mathcal{D}$. Hence, $F^{\prime}\prec F$ by Definition 3.5 (i). Thus, we take $\tilde{F}=F^{\prime}$.
Now suppose that $s_{1}=\alpha_{s}+4$. First, let $\mu\neq\alpha_{s}+2$. We have $13\leq\alpha_{s}<\alpha_{s}+4=s_{1}<n-$ 1. It follows that $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. Moreover, $\alpha_{s}\nsim s_{1}$, and hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.6, and we take $\tilde{F}=F^{\prime}$. So assume that $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime}\in\mathcal{T}_{s}$ and $F^{\prime}\ll F$ by Proposition 3.12. Since $\alpha_{s}<m-1$, we have $F,F^{\prime}\notin\mathcal{D}$. Therefore, $F^{\prime}\prec F$ by Definition 3.5 (i), and we take $\tilde{F}=F^{\prime}$.
Now, let $\mu=\alpha_{s}+2$. Let $F^{\prime\prime c}:=(F^{c}\setminus\{s_{1}\})\sqcup\{\mu\}=\{\alpha_{s},\mu,n-1\}$. Since $13\leq\alpha_{s}<\alpha_{s}+2=\mu=s_{1}-2<s_{1}<n-$ 1, it follows that $\alpha_{s}\nsim n-1$ and $\mu\nsim n-1$. Therefore $F^{\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime\prime}\prec F$ by Proposition 3.6 and we take $\tilde{F}=F^{\prime\prime}$. So, let $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime\prime}\in\mathcal{T}_{s}$ and thus, $F^{\prime\prime c}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. Since $\mu=\alpha_{s}+2<\alpha_{s}+4=s_{1}$, we have $F^{\prime\prime}\ll F$ by Definition 3.4 (i). Also, $\alpha_{s}<m-1$ implies that $F,F^{\prime\prime}\notin\mathcal{D}$. Hence, $F^{\prime\prime}\prec F$ by Definition 3.5 (i) and we take $\tilde{F}=F^{\prime\prime}$.
Therefore, in each case, we have a facet $\tilde{F}\in M(\Delta_{3}(C_{n}^{2}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$. Thus, $F$ is a spanning facet. ∎
@@
Since $F\in\mathcal{S}$, we have $F\in\mathcal{T}_{s}$ for some $s$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$. Let $\mathcal{U}=V(C_{n}^{2})\setminus\bigl{(}\{\alpha_{s}-4\ (\text{mod $n$}),\alpha_{s}-3,\alpha_{s}-2,\ldots,2m-\alpha_{s}-1\}\cup\{n-1\}\bigr{)}$.
@@
    We have $\alpha_{s}=3$ and $s_{1}\in V(C_{n}^{2})\setminus(\{0,1,2,3,\ldots,2m-4\}\sqcup\{n-1\})$. Observe that $s_{1}\in\mathcal{U}$. Therefore, $F$ is a spanning facet by Proposition 3.13.
@@
    We have $\alpha_{s}\in\{4,5,\ldots,m-2\}$ and $s_{1}\in V(C_{n}^{2})\setminus\{\alpha_{s}-4,\alpha_{s}-3,\alpha_{s}-2\}\sqcup\{\alpha_{s},\alpha_{s}+1,\ldots,2m-\alpha_{s}-1\}\sqcup\{n-1\}$.
@@
    So assume that $s_{1}=\alpha_{s}-1$. We have $m\leq n-4$ (as $n\geq 9$). We consider three cases: $\mu<\alpha_{s}$, $\mu\in\{\alpha_{s}+1,\alpha_{s}+2\}$, and $\mu>\alpha_{s}+2$.
    First, let $\mu<\alpha_{s}$. Then $\mu<s_{1}=\alpha_{s}-1$ (as $\mu\in F$). Moreover, $\alpha_{s}\leq m-2$ implies that $\alpha_{s}<_{\mathcal{O}}\mu$. Hence $F^{\prime\prime}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. Since $64\leq\alpha_{s}\leq m-2\leq n-$ 6, we have $\alpha_{s}\nsim n-1$. Further, $\mu<\alpha_{s}$ implies that either $\mu\in\{0,1\}$, or $2\leq\mu<\alpha_{s}$. Hence, $\mu\nsim\alpha_{s}$ or $\mu\nsim n-1$. Therefore $F^{\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $\mu<s_{1}$, we have $F^{\prime\prime}\ll F$ by Definition 3.4 (i). Moreover, $\alpha_{s}<m-1$ implies that $F,F^{\prime\prime}\notin\mathcal{D}$. Therefore, $F^{\prime\prime}\prec F$ Definition 3.5 (i) and we take $\tilde{F}=F^{\prime\prime}$.
    Now, let $\mu\in\{\alpha_{s}+1,\alpha_{s}+2\}$. We have $14\leq\alpha_{s}<\mu\leq\alpha_{s}+2\leq m\leq n-4<n-$ 1. Thus, $\alpha_{s}\nsim n-1$ and $\mu\nsim n-1$. Hence $F^{\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $\alpha_{s}\leq m-2$, we have $\mu<_{\mathcal{O}}\alpha_{s}$. Therefore, $F^{\prime\prime}\prec F$ by Proposition 3.6, and we take $\tilde{F}=F^{\prime\prime}$.
    Lastly, assume that $\mu>\alpha_{s}+2$. Then $13\leq\alpha_{s}-1=s_{1}<\alpha_{s}<\alpha_{s}+2<\mu<n-$ 1 implies that $\mu\nsim\alpha_{s}$ and $\mu\nsim s_{1}$. Hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.6, and we take $\tilde{F}=F^{\prime}$. Now, if $\alpha_{s}<_{\mathcal{O}}\mu$, then $F^{\prime}\in\mathcal{T}_{s}$ and $F^{\prime}\ll F$ by Proposition 3.12. Since $\alpha_{s}<m-1$, we have $F,F^{\prime}\notin\mathcal{D}$. Hence, $F^{\prime}\prec F$ by Definition 3.5 (i). Therefore, we take $\tilde{F}=F^{\prime}$.
@@
    We have $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},n-1\}$, where $\alpha_{s}\in\{m-1,m,m+1,m+2,\ldots,n-3\}$, $s_{1}\in V(C_{n}^{2})\setminus(\{\omega,\omega+1,\ldots,\alpha_{s},\alpha_{s}+1,\alpha_{s}+2,\alpha_{s}+3\text{ (mod $n$})\}\sqcup\{y\})$, $\omega=\min\{2m-\alpha_{s},\alpha_{s}-4\}$ and
@@
    We have $m\geq 5$ (as $n\geq 9$). First, suppose that $s_{1}\neq\alpha_{s}+4$ (mod $n$). Since $s_{1}\notin\{\omega,\omega+1,\ldots,\alpha_{s}+3\text{ (mod $n$)}\}$ and $\omega\leq\alpha_{s}-4$, it follows that $s_{1}<\alpha_{s}-4$ or $s_{1}>\alpha_{s}+4$ (mod $n$). Thus $|\alpha_{s}-s_{1}|\geq 5$. Now, since $\alpha_{s}\geq m-1\geq 4$ and $s_{1}<n-1$, we get $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. We have $\alpha_{s}\nsim s_{1}$. Therefore $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.6, and we take $\tilde{F}=F^{\prime}$. Now, if $\alpha_{s}<_{\mathcal{O}}\mu$, then $F^{\prime}\in\mathcal{T}_{s}$ and $F^{\prime}\ll F$ by Proposition 3.12. Since $s_{1}<\alpha_{s}-4$ or $s_{1}>\alpha_{s}+4$ (mod $n$), we have $F,F^{\prime}\notin\mathcal{D}$. Hence $F^{\prime}\prec F$ by Definition 3.5 (i). Therefore, we take $\tilde{F}=F^{\prime}$.
    Now, let $s_{1}=\alpha_{s}+4$ (mod $n$). Observe that if $\alpha_{s}\geq n-5$, then $\alpha_{s}+4\ \text{(mod $n$)}=y$. Therefore, since $s_{1}\neq y$, we have $\alpha_{s}\leq n-6$. This means that $\alpha_{s}+4\ \text{(mod $n$)}=\alpha_{s}+4$. First, let $\mu\notin\{\alpha_{s}+2,\alpha_{s}+3\}$. We have $14\leq m-1\leq\alpha_{s}<\alpha_{s}+4=s_{1}<n-$ 1. It follows that $\mu\nsim\alpha_{s}$ or $\mu\nsim s_{1}$. Further, since $\alpha_{s}\nsim s_{1}$, we get $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$.
    If $\mu<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$ by Proposition 3.6, and we take $\tilde{F}=F^{\prime}$. So assume that $\alpha_{s}<_{\mathcal{O}}\mu$. Then $F^{\prime}\in\mathcal{T}_{s}$ and $F^{\prime}\ll F$ by Proposition 3.12. Since $n-1>s_{1}=\alpha_{s}+4$ and $\mu\neq\alpha_{s}+3$, we have $F,F^{\prime}\notin\mathcal{D}$. Therefore, $F^{\prime}\prec F$ by Definition 3.5 (i), and we take $\tilde{F}=F^{\prime}$.
    Now, let $\mu\in\{\alpha_{s}+2,\alpha_{s}+3\}$. Observe that since $\alpha_{s}\geq m-1$, we have $\alpha_{s}<_{\mathcal{O}}\mu$, and hence $F^{\prime\prime}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. We have $14\leq m-1\leq\alpha_{s}<\alpha_{s}+2<\alpha_{s}+3\leq n-3<n-$ 1 (as $\alpha_{s}\leq n-6$). Clearly, $\alpha_{s}\nsim n-1$. Moreover, if $\mu=\alpha_{s}+2$, then $\mu\nsim n-1$; and if $\mu=\alpha_{s}+3$, then $\mu\nsim\alpha_{s}$. Thus $F^{\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $\mu<s_{1}$, we have $F^{\prime\prime}\ll F$. Further, $\alpha_{s}\leq n-6$ implies that $n-1\geq\alpha_{s}+5$ and thus, $F,F^{\prime\prime}\notin\mathcal{D}$. Hence, $F^{\prime\prime}\prec F$ by Definition 3.5 (i), and we take $\tilde{F}=F^{\prime\prime}$.
    Thus, in each case, we have a facet $\tilde{F}\in M(\Delta_{3}(C_{n}^{2}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$. Thus, $F$ is a spanning facet.
@@
Suppose $F\in M(\Delta_{3}(C_{n}^{2}))$ be such that either $F^{c}=\{x,x+1\ (\text{mod $n$}),x+4\ (\text{mod $n$})\}$, or $F^{c}=\{x,x+3\ (\text{mod $n$}),x+4\ (\text{mod $n$})\}$ for some $x\in V(C_{n}^{2})$. Then $F$ is not a spanning facet.
@@
We have $x+2$ (mod $n$) $\in F$. Let ${F^{\nu}}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{x+2\text{ (mod $n$)}\}$ for $\nu\in F^{c}$. Then for any $\nu\in F^{c}$, $C_{n}^{2}[{F^{\nu}}^{c}]$ is connected. Hence, there does not exist a facet $\tilde{F}$ such that $\tilde{F}\prec F$ and $\tilde{F}\cap F=F\setminus\{x+2\text{ (mod \ $n$)}\}$. Therefore, $F$ is not a spanning facet by definition. ∎
@@
Observe that if $F\in\mathcal{D}$, then either $F^{c}=\{x,x+1\ (\text{mod \ $n$}),x+4\ (\text{mod \ $n$})\}$ or $F^{c}=\{x,x+3\ (\text{mod \ $n$}),x+4\ (\text{mod \ $n$})\}$ for some $x\in V(C_{n}^{2})$. Therefore, if $F\in\mathcal{D}$, then $F$ is not a spanning facet by Proposition 3.15.
@@
Let $F,F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Let $F\in\mathcal{T}_{s}$ and $F^{\prime}\in\mathcal{T}_{t}$ for some $s\neq t$ such that ${F}^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$ and ${F^{\prime}}^{c}=\{\mu,s_{1},s_{2}\}$ for some $\mu\in F$. If $\alpha_{s}<_{\mathcal{O}}\mu$, then $F\prec F^{\prime}$.
@@
Let $F\in M(\Delta_{3}(C_{n}^{2}))$ be a spanning facet such that $F\in\mathcal{T}_{s}$ for some $s$ and $F^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$. Then $\alpha_{s}\notin\{0,1,2,n-2,n-1\}$ and $s_{2}=n-1$.
@@
By Remark 3.3, $\alpha_{s}\notin\{0,1,n-1\}$. We have $C_{n}^{2}[F^{c}]$ is disconnected and by Remark 3.2, $\alpha_{s}<_{\mathcal{O}}s_{1},s_{2}$. Therefore, since $F$ is a spanning facet, we get $\alpha_{s}\notin\{2,n-2\}$ by Proposition 3.15. Hence $\alpha_{s}\notin\{0,1,2,n-2,n-1\}$.
Observe that $\alpha_{s}<_{\mathcal{O}}n-1$. Suppose $n-1\notin F^{c}$. Then Remark 3.11 implies that there exists $\tilde{F}\in M(\Delta_{3}(C_{n}^{2}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{n-1\}$ for some $\nu\in F^{c}$.